# Can We Benefit from Reconfigurable Intelligent Surfaces in Upper Mid-Band 6G Networks? A Critical Look for Promising Use Cases

### v1

@@ Abstract
Firstly, the placement of large RIS can be a showstopper for MNOs, especially in densely populated urban areas. It is not only difficult to find suitable deployment locations, but it also likely incurs high costs due to site rentals. MNOs may not find such deployments financially viable unless an RIS can be hidden behind other equipment such as commercial billboards.
There are also technical challenges related to having many RIS elements. In this paper, we have assumed that perfect channel state information (CSI) is always available. However, estimating 10 000 channel coefficients is challenging, not only from a signal processing perspective but also in terms of the immense pilot load. The latter problem can potentially be addressed by exploiting known channel characteristics, such as the existence of LoS paths whose angles can be estimated [14]. The use of such algorithms in practical environments with hardware imperfections and multipath components is yet to be demonstrated. Furthermore, our results were obtained using the optimal phase-shifts, which require infinite-resolution controllers at the RIS. Although near-optimal performance can be achieved with lower resolution, implementing a control unit for huge RIS element numbers can be very hard and costly.
An RIS has the ability to manipulate signals and redirect them in desired directions for MNOs. However, after reflecting off the RIS, the RF waves will again reflect from the wireless medium and scatterers, eventually causing interference. Additionally, in commercial systems, multiple MNOs operate in adjacent bands, and current RIS technology cannot filter out only the band of a single MNO but will affect all MNOs. Hence, if one MNO deploys an RIS and operates it according to one of the aforementioned use cases, the channel conditions in the adjacent MNOs’ bands will also be modified. This leads to extra small-scale fading in the best case and inter-operator pilot contamination in the worst case [15]. This raises questions about whether each operator can have its own RISs or if all RISs must be shared and jointly operated by all MNOs. Since the upper mid-band is also adjacent to other sensitive systems, such as satellite links, the impact of RIS on coexistence must be carefully studied.
@@ Conclusions
In this paper, we have studied 6G upper mid-band systems and identified four prospective use cases for RIS deployments. Enhancing the range and capacity for FWA has a clear commercial value since more customers can be accommodated per base station. There are multiple ways to deploy RISs to improve the SE of mobile users, but when considering realistic channel models with a non-zero static path and uncertain LoS to the RIS, the gains become rather modest. The required physical RIS size can also be a showstopper. However, an RIS can greatly enhance network reliability for cell-edge users by combating fading through massive channel hardening. This might be the most promising use case. Many practical problems must be solved before any commercial RIS deployments can begin, including the coexistence between MNOs and other services.

### v2

@@ Abstract
Deployment of RIS and Site Rental Costs: Firstly, the placement of large RIS can be a showstopper for MNOs, especially in densely populated urban areas. It is not only difficult to find suitable deployment locations, but it also likely incurs high costs due to site rentals. MNOs may not find such deployments financially viable unless an RIS can be hidden behind other equipment such as commercial billboards.
Control of RIS and Interference Management: An RIS has the ability to manipulate signals and redirect them in desired directions for MNOs. However, after reflecting off the RIS, the RF waves will again reflect from the wireless medium and scatterers, eventually causing interference. Additionally, in commercial systems, multiple MNOs operate in adjacent bands, and current RIS technology cannot filter out only the band of a single MNO but will affect all MNOs. Hence, if one MNO deploys an RIS and operates it according to one of the aforementioned use cases, the channel conditions in the adjacent MNOs’ bands will also be modified. This leads to extra small-scale fading in the best case and inter-operator pilot contamination in the worst case [13]. This raises questions about whether each operator can have its own RISs or if all RISs must be shared and jointly operated by all MNOs. Since the upper mid-band is also adjacent to other sensitive systems, such as satellite links, the impact of RIS on coexistence must be carefully studied.
Channel Estimation and Phase-Shift Optimization with Large RIS: In this paper, we have assumed that perfect channel state information (CSI) is always available. However, estimating 10 000 channel coefficients is challenging, not only from a signal processing perspective but also in terms of the immense pilot load. The latter problem can potentially be addressed by exploiting known channel characteristics, such as the existence of LoS paths whose angles can be estimated [14]. The use of such algorithms in practical environments with hardware imperfections and multipath components is yet to be demonstrated. Furthermore, our results were obtained using the optimal phase shifts, which require infinite-resolution controllers at the RIS. Although near-optimal performance can be achieved with lower resolution, implementing a control unit for huge RIS element numbers can be very hard and costly.
Near-Field Effects: In this paper, we adopted the plane-wave assumption, assuming all nodes are located in each others’ far-field regions. However, with large RIS sizes in the upper-mid band, radiative near-field effects will be inevitable when the users are close to the RIS (e.g., in Use Cases 3 and 4). In these situations, we end up with spherical waves with non-linear amplitude and phase variations across the RIS [10]. The simulation results in this paper can be seen as upper bounds since the radiative near-field effects can lead to power losses. Nevertheless, as shown in [15], with proper beamfocusing/phase optimization, this performance loss can be minimized. The near-field effect can also increase the MIMO rank, which can improve the performance. It is essential to develop an industry-accepted near-field-compliant channel model for RIS-assisted communication so performance can be properly evaluated and algorithms designed to capture all essential effects, there are more yet to be explored.
@@ Conclusions
In this paper, we have studied RIS deployments in the upper mid-band to answer whether RIS can provide substantial gains or mostly cause technical and operational complexity (i.e., pain) with a marginal gain. First, we have demonstrated that a LoS link is crucial for RIS deployments in upper mid-band systems. Without it, RIS might be painful by introducing the need for highly complex algorithms (i.e., channel estimation and phase-shift design for large elements) in return for a subtle gain. Based on this observation, we have identified four prospective use cases for RIS deployments where an RIS node significantly improves performance. Within these four use cases, enhancing the range and capacity for FWA has a clear commercial value since more customers can be accommodated per base station. There are multiple ways to deploy RISs to improve the SE of mobile users, but when considering realistic channel models with a non-zero static path and uncertain LoS to the RIS, the gains become rather modest. The required physical RIS size can also be a showstopper. Nevertheless, an RIS can greatly improve performance, particularly when located near the end nodes. This could either refer to a blind spot with multiple users or a cell-edge user. In these cases, RIS-assisted communication guarantees a substantial increase in SE for a blind spot or greatly enhanced network reliability for cell-edge users by combating fading through channel hardening. These might be the most promising use cases. Many practical problems must be solved before any commercial RIS deployments can begin, including the coexistence between MNOs and other services.